# AFABench: A Generic Framework for Benchmarking Active Feature Acquisition

### v1

@@ A.1.6. OP-MFRL
  • We use the same architecture for (Fashion)MNIST as originally presented: $[512,512,128,64]$. For the remaining datasets, we use the architecture originally used in the diabetes dataset, $[64,32,16]$.
@@
  • We do not use an adaptive learning rate.
  • Original OL originally includes the stop action, although the method is designed for the hard-budget case. We remove the stop action.
  • The OL repo only provides the number of training episodes for the diabetes dataset. As for all other RL methods, we pretrain until convergence and train with a consistent number of batches.
  • Original OL updates the target network in discrete steps, but we use a soft update instead, with $\tau=0.005$.
  • Original OL trains a single agent with batch size 1. We tried this, but it made the method very unstable. Consequently, we increased the number of agents and batch size to be consistent across all RL methods.
  • Original OL decays $\epsilon$ exponentially, but we decrease it linearly during the first half of training. This makes it consistent with our JAFA implementation.
  • Original OL prescribes a replay buffer of size $1000n_{\textrm{features}}$. We use this for all datasets except (Fashion)MNIST where the larger samples required a decreased buffer size of $30000$ to fit everything on the GPU.
@@
  • This is by far the fastest of the RL methods, due to the simple architecture.
  • However, this simplicity comes at a cost: the method cannot differentiate between unobserved features and observed features with value 0. This is a significant drawback that makes the method perform especially poorly on datasets with a large number of zeros.
A.1.7. AACO
@@
  • The AACO method cannot choose a subset of features in a data instance; rather, it will break ties with selecting the feature that minimizes the loss, as per the original implementation.
  • The method cannot terminate early, as our benchmark requires a hard budget.
@@
  • To adhere to the fair benchmark, the original classifiers suggested by Valancius et al. were not used. Instead, the shared classifier was used.
A.1.8. Common across Discriminative Greedy Methods
@@
For both discriminative greedy methods, we use the same predictor structure to ensure fair comparison. Specifically, we use a predictor with two hidden layers of size [128,128] and a dropout rate of 0.3. Both methods involve predictor pretraining and joint training of the predictor and selector/estimator. During training, we concatenate the masked features with the feature mask as the model input according to the default settings of (Covert et al., 2023; Gadgil et al., 2024). We use a batch size of 128 and a learning rate of 0.001 during both model pretraining and training, and we use early stopping in both cases.
A.1.9. GDFS-DG
@@
We use the default gradual temperature annealing strategy (Covert et al., 2023) to train the models during the joint training phase. During model training, we gradually decrease the temperature from 1.0 to 0.1, with a total of five temperature stages using logarithmic spacing. At each temperature stage, we train the model for a maximum of 250 epochs. We use the model with the lowest validation loss at each temperature as the starting point for the next temperature stage and will continue to train until the best model is obtained in the final stage.
A.1.10. DIME-DG
@@
During the joint training phase of DIME-DG, we perform feature selection using the $\epsilon$-greedy strategy with an initial $\epsilon$ set to 0.05. If the validation accuracy does not improve over five consecutive epochs, we reduce the $\epsilon$ by multiplying it by 0.2, up to a maximum of ten times. We also use the default bounded strategy (Gadgil et al., 2024) for CMI estimation. We jointly train the model for a maximum of 250 epochs.
A.1.11. EDDI-GG
After joint pretraining of partial variational autoencoder and predictor, the EDDI-GG model does not require further training. Hence, we only save the same models for different budget values in the training stage and use this model for subsequent evaluation.
A.1.12. Common across Static Methods
@@
During the training phase of the static method, we use a selector and classifier with two hidden layers of size [128,128] for both methods. We train both methods with a batch size of 128 and a learning rate of 0.001. We train each method for a maximum of 250 epochs with an early stopping strategy.
A.1.13. CAE-S
To ensure fair comparison,instead of selecting the optimal features for each budget value from 1 to the maximum feature budget, we use the CAE selector to only select the optimal features for the maximum budget. We then use this selected feature list to train a separate predictor for each budget.
A.2. Hardware
All experiments were performed on NVIDIA T4 GPUs.
@@ A.2. Hardware
A.3. Proof that $\lambda=0.75$ gives maximum weight to 4-step return
@@
A.4. Additional Results
This section supplements the experimental results in different budget settings, covering the three budget ranges. We evaluate the performance of different methods using both the builtin classifiers and the shared classifiers, where the results are illustrated in Figures˜A.1, A.2, A.3, A.4 and A.5. These experiments are mainly conducted to evaluate the robustness of our conclusions in different experimental settings.
From the experimental results, we observe that the overall trends are consistent with the findings in the main paper under the large budget setting, with the relative performance rankings of each method mostly remaining stable. We also observe that in most cases the results using builtin classifiers are slightly better than those using shared classifiers. This outcome is expected, as the builtin classifier is usually jointly optimized with the feature selection strategy.
We also observe that with a small budget, reinforcement learning methods tend to select more informative features earlier, resulting in behaviors that more closely resemble greedy methods. In contrast, for large budgets, their selection strategy is relatively more balanced. This observation confirms our initial hypothesis that the behavior of reinforcement learning methods is affected by budget constraints. As an example, the JAFA-MFRL method achieves $\approx 55\%$ accuracy with budget $10$ on FashionMNIST when evaluated with the shared classifier (see Figure 1(d)). In contrast, the same method only obtains $\approx 45\%$ accuracy after $10$ features when trained with budget $30$ on the same dataset (see Figure 1(d)).
(a) AFAContext (Feature Budget = 5)
(b) cube (Feature Budget = 5)
(c) diabetes (Feature Budget = 10)
(d) FashionMNIST (Feature Budget = 20)
(e) physionet (Feature Budget = 10)
(f) MNIST (Feature Budget = 20)
(g) miniboone (Feature Budget = 10)
Figure A.1. Performance of Different Methods on Seven Datasets Using Shared Classifiers (Medium Budgets)
(a) AFAContext (Feature Budget = 3)
(b) cube (Feature Budget = 3)
(c) diabetes (Feature Budget = 5)
(d) FashionMNIST (Feature Budget = 10)
(e) physionet (Feature Budget = 5)
(f) MNIST (Feature Budget = 10)
(g) miniboone (Feature Budget = 5)
Figure A.2. Performance of Different Methods on Seven Datasets Using Shared Classifiers (Small Budgets)
(a) AFAContext (Feature Budget = 10)
(b) cube (Feature Budget = 10)
(c) diabetes (Feature Budget = 15)
(d) FashionMNIST (Feature Budget = 30)
(e) physionet (Feature Budget = 15)
(f) MNIST (Feature Budget = 30)
(g) miniboone (Feature Budget = 15)
Figure A.3. Performance of Different Methods on Seven Datasets Using Builtin Classifiers (Large Budgets)
(a) AFAContext (Feature Budget = 5)
(b) cube (Feature Budget = 5)
(c) diabetes (Feature Budget = 10)
(d) FashionMNIST (Feature Budget = 20)
(e) physionet (Feature Budget = 10)
(f) MNIST (Feature Budget = 20)
(g) miniboone (Feature Budget = 10)
Figure A.4. Performance of Different Methods on Seven Datasets Using Builtin Classifiers (Medium Budgets)
(a) AFAContext (Feature Budget = 3)
(b) cube (Feature Budget = 3)
(c) diabetes (Feature Budget = 5)
(d) FashionMNIST (Feature Budget = 10)
(e) physionet (Feature Budget = 5)
(f) MNIST (Feature Budget = 10)
(g) miniboone (Feature Budget = 5)
Figure A.5. Performance of Different Methods on Seven Datasets Using Builtin Classifiers (Small Budgets)

### v3

@@ E.1.8. OL-MFRL+no-mask
  • We used the same architecture for (Fashion)MNIST as originally presented: $[512,512,128,64]$. For the remaining datasets, we used the architecture originally used for the Diabetes dataset, $[64,32,16]$.
@@
  • We did not use an adaptive learning rate.
  • The OL repo only provides the number of training episodes for the Diabetes dataset. As for all other RL methods, we pretrained until convergence and trained with a consistent number of batches.
  • Original OL updated the target network in discrete steps, but we used a soft update instead, with $\tau=0.005$.
  • Original OL trained a single agent with batch size 1. We tried this, but it made the method very unstable. Consequently, we increased the number of agents and batch size to be consistent across all RL methods.
  • Original OL decayed $\epsilon$ exponentially, but we decreased it linearly during the first half of training. This makes it consistent with our JAFA implementation.
  • Original OL prescribed a replay buffer of size $1000n_{\textrm{features}}$. We used a fixed size of 100 000100\,000 instead.
@@
  • The architectural simplicity comes at a cost: the method cannot differentiate between unobserved features and observed features with value 0. This is a significant drawback that makes the method perform especially poorly on datasets with a large number of zeros. Due to this deficiency, we also investigated a variant of OL which concatenated the feature mask to the input.
E.1.9. OL-MFRL
OL-MFRL makes no additional modifications to OL-MFRL+no-mask apart from concatenating the feature mask to the input. This changes the size of the first layers in both the P-network and Q-network.
E.1.10. AACO
@@
  • The AACO method cannot execute an entire selected subset in one step; instead, it breaks ties by selecting the feature that minimizes expected loss (without an explicit cost term), as in the original implementation.
  • In the hard-budget setting, the method cannot terminate early.
@@
  • To adhere to the fair benchmark, the original classifiers suggested by Valancius et al. (i.e., XGBClassifier and Bayes classifier) were not used. Instead, the external classifier was used.
E.1.11. AACO+NN
For AACO+NN, we first loaded a trained AACO object and generated oracle rollouts on the training split. Rollouts used the same Initializer and Unmasker as evaluation. In soft-budget runs, the oracle could stop early; in hard-budget runs, forced acquisition was enabled and the rollout length was capped by the budget.
The policy network was a masked-MLP-style model taking concatenated masked features and feature mask as input, with hidden layers $[256,256]$, dropout $0.1$, and an output dimension equal to the number of selections plus one stop action. We trained with AdamW (learning rate $10^{-3}$), batch size $256$, 10% validation split, and early stopping with patience 10 (maximum 100 epochs). The final saved method combines this policy with the same classifier object used for AACO predictions.
E.1.12. Common across Discriminative Myopic Methods
@@
For both discriminative myopic methods, we used the same predictor setup for each modality to ensure a fair comparison. Specifically, for tabular datasets, we used a predictor with two hidden layers of size [128,128] and a dropout rate of 0.3, whereas for Imagenette we used a pretrained ResNet-50 (He et al., 2015) predictor. Both methods involved predictor pretraining and joint training of the predictor and selector/estimator. During training, for tabular datasets, we concatenated the masked features with the corresponding feature mask as the model input according to the default settings of (Covert et al., 2023; Gadgil et al., 2024), while for Imagenette we used the masked images directly as the model input. We used a batch size of 128 during both model pretraining and training, and we used early stopping in both cases.
E.1.13. GDFS-DM
@@
We used the default gradual temperature annealing strategy (Covert et al., 2023) to train the models during the joint training phase. For tabular datasets, we used an initial learning rate of $1\times 10^{-3}$ and gradually decreased the temperature from 1.0 to 0.1, with a total of five temperature stages using logarithmic spacing. At each temperature stage, we trained the model for a maximum of 250 epochs. We used the model with the lowest validation loss at each temperature as the starting point for the next temperature stage. The checkpoint with the lowest validation loss across all stages was selected as the final model. For Imagenette, we used the same annealing schedule, but with an initial learning rate of $1\times 10^{-5}$, a minimum learning rate of $1\times 10^{-8}$, and a maximum of 30 epochs per stage.
E.1.14. DIME-DM
@@
During the joint training phase of DIME-DM, we performed feature selection using the $\epsilon$-greedy strategy with the initial $\epsilon$ set to 0.05. For tabular datasets, if the validation accuracy did not improve over five consecutive epochs, we reduced the $\epsilon$ by multiplying it by 0.2, up to a maximum of ten times. We also used the default bounded strategy (Gadgil et al., 2024) for CMI estimation. We jointly trained the model for a maximum of 250 epochs. For Imagenette, we used the same $\epsilon$-greedy strategy with the same initial $\epsilon$ and decay factor, but reduced $\epsilon$ after three consecutive epochs without validation accuracy improvement and trained the model for up to 150 epochs. For joint training on tabular datasets, we tied the first two linear layers of the predictor and estimator so that they share the same module. For joint training on Imagenette, we used a shared backbone for both the predictor and the estimator.
E.1.15. EDDI-GM
After joint pretraining of the PVAE and predictor, the EDDI-GM model did not require further training. Hence, we only saved the same models for different budget values and classifier settings in the training stage and used this model for subsequent evaluation. During evaluation, we considered two variants of the EDDI method in this paper. They shared the same feature selection procedure, but differed in how predictions were computed during feature selection. In the EDDI-GM+builtin variant, the method performed multiple Monte Carlo samplings based on the current masked features to obtain latent representations, and used the PVAE’s built-in classifier to compute the current predictions by averaging over these representations. The same Monte Carlo samples were then used to complete missing features, and candidate inputs were constructed by sequentially revealing each input feature. Predictions were recomputed on these candidate inputs and compared with the original predictions to measure the informativeness of each feature. In contrast, in the EDDI-GM variant, the original predictions were made directly based on the masked features, and the sampler was only used to complete missing feature values. During feature informativeness evaluation, predictions were computed directly on the completed candidate inputs without resampling latent representations. We used a total of 128 Monte Carlo samples in both variants.
E.1.16. Common across Static Methods
@@
During the training phase of the static methods on tabular datasets, we used a selector and classifier with two hidden layers of size [128,128] for both methods. We trained both methods with a batch size of 128 and a learning rate of 0.001.
E.1.17. CAE-S
For tabular datasets experiments, we trained the selector for up to 2500 epochs and the predictor for up to 250 epochs, both with an early stopping strategy. For Imagenette experiments, we used a pretrained ResNet-50 (He et al., 2015) for both the selector and classifier. The selector was trained for up to 500 epochs and the classifier was trained for up to 50 epochs. Both the selector and the classifier were trained with a learning rate of $1\times 10^{-5}$. For the selector training on both tabular and image datasets, we used 10 Gumbel-Softmax temperature annealing steps.
E.1.18. PT-S
For tabular dataset experiments, we trained both the selector and the classifier for a maximum of 250 epochs with an early stopping strategy.
E.2. Hardware
  • Myopic methods on the Imagenette dataset used NVIDIA A40 GPUs during the training stage.
  • AACO(+NN) used 4 AMD EPYC 9354 Zen4 CPU cores during both training and evaluation.
  • The remaining methods used one NVIDIA T4 GPU during both (pre)training and evaluation.
@@ E.2. Hardware
E.3. Proof that $\lambda=0.75$ gives maximum weight to 4-step return
@@
Appendix F Proof of Theorem 4.1
We begin by stating the theorem more formally. Note that we also include the worst-case and best-case cost here.
Theorem F.1 (Noiseless CUBE-NM query complexity).
Consider the CUBE-NM dataset with $n_{c}$ contexts in the noiseless setting $\sigma=0$, so that each instance $x\in\mathcal{X}$ determines its label with full certainty (i.e., $H(\mathbf{y}\mid x)=0$). Let $f(S,x_{S})=p(\mathbf{y}\mid x_{S})$ be Bayes optimal and assume uniform feature costs $c_{i}=1$ for all $i\in[d]$. In addition, assume we allow any policy to acquire features until $H(\mathbf{y}\mid x_{S})=0$, and denote $\Pi$ as the set of all such policies. Let $\pi^{\ast}=\operatorname*{argmin}_{\pi\in\Pi}\mathbb{E}_{\mathbf{x}}\mathbb{E}_{\pi}[c(\pi[\mathbf{x}])]$ denote an optimal policy in terms of expected cost and let $\pi_{\mathrm{CMI}}\in\Pi$ denote the myopic CMI policy in (3). Then $\mathbb{E}_{\mathbf{x}}\mathbb{E}_{\pi^{\ast}}[c(\pi^{\ast}[\mathbf{x}])]=\mathbf{1}\{n_{c}\geq 2\}+2.25$, with $\min_{x\in\mathcal{X}}c(\pi^{\ast}[x])=\mathbf{1}\{n_{c}\geq 2\}+1$ and $\max_{x\in\mathcal{X}}c(\pi^{\ast}[x])=\mathbf{1}\{n_{c}\geq 2\}+3$. In contrast, $\mathbb{E}_{\mathbf{x}}\mathbb{E}_{\pi_{\mathrm{CMI}}}[c(\pi_{\mathrm{CMI}}[\mathbf{x}])]=13(2n_{c}+1)/16$, with $\min_{x\in\mathcal{X}}c(\pi_{\mathrm{CMI}}[x])=1$ and $\max_{x\in\mathcal{X}}c(\pi_{\mathrm{CMI}}[x])=3n_{c}$.
Proof.
Let $n=n_{c}$. In the noiseless setting $\sigma=0$, each label-context pair corresponds to a deterministic feature vector.
Notation. For noiseless CUBE, let $x^{(y)}\in\{0,0.5,1\}^{10}$ denote the 10-dimensional prototype for class $y\in\{1,\dots,8\}$ (restricted to the first 10 coordinates), and let $x^{(y)}_{i}$ be its $i$-th coordinate. For CUBE-NM with $n$ contexts, each instance has one context feature $f_{1}$ with value $x_{f_{1}}\in\{1,\dots,n\}$, and $n$ disjoint 10-dimensional blocks. Write $g_{b,i}$ for the $i$-th feature inside block $b\in\{1,\dots,n\}$. If the label is $y$ and the active block is $j$, then $x_{g_{j,i}}=x^{(y)}_{i}$ for all $i\in\{1,\dots,10\}$ and $x_{g_{b,i}}=0.5$ for all $b\neq j$. Unit costs mean $c(\pi[x])$ equals the number of queried features.
Step 1: noiseless CUBE costs. Consider the noiseless CUBE core on 10 features, with 8 classes and deterministic prototypes $x^{(y)}\in\{0,0.5,1\}^{10}$ for $y\in\{1,\dots,8\}$. A policy adaptively queries feature indices $i\in\{1,\dots,10\}$ and observes $x^{(y)}_{i}$.
An upper bound (a policy with expected cost $2.25$). Query feature $7$. If $x^{(y)}_{7}=1$, then $y=5$ and stop. If $x^{(y)}_{7}=0$, then $y\in\{6,7\}$ and querying feature $6$ separates them. If $x^{(y)}_{7}=0.5$, then $y\in\{1,2,3,4,8\}$; query feature $2$, which isolates $y=1$ and $y=2$, and leaves $y\in\{3,4,8\}$, which is resolved by one additional query. Under the uniform class prior, the expected number of queries is $1+(2/8)\cdot 1+(5/8)\cdot(1+(3/5)\cdot 1)=2.25$. The best case is $1$ query (when $x^{(y)}_{7}=1$), and the worst case is $3$ queries.
Optimality (dynamic programming on belief states). A noiseless state is fully described by a pair $(B,S)$, where $B\subseteq\{1,\dots,8\}$ is the set of labels still consistent with the observed feature values, and $S\subseteq\{1,\dots,10\}$ is the set of already queried features. Let $V(B,S)$ denote the minimum expected number of additional queries needed to identify the label when the queried set is $S$ and remaining labels are $B$ (note that the outcomes of the features are irrelevant, because $(B,S)$ is a sufficient statistic at any state for this simple dataset, see (Aronsson et al., 2026) for more on this). The boundary condition is $V(\{y\},S)=0$ for all $y$ and $S$.
If a feature $i\notin S$ is queried next and the observed value is $v\in\{0,0.5,1\}$, the remaining possible labels become $B_{i=v}=\{y\in B:x^{(y)}_{i}=v\}$ and the queried set becomes $S\cup\{i\}$. In the noiseless uniform setting, $\Pr(v\mid B,S,i)=|B_{i=v}|/|B|$. Therefore, for $|B|\geq 2$,
$V(B,S)=1+\min_{i\in\{1,\dots,10\}\setminus S}\sum_{v\in\{0,0.5,1\}}\Pr(v\mid B,S,i)\,V(B_{i=v},S\cup\{i\}).$
Evaluating this recursion on the noiseless CUBE dataset yields $V(\{1,\dots,8\},\emptyset)=2.25$ (note that this is straightforward due to the simplicity of the dataset and $\sigma=0$). Since the policy above attains expected cost $2.25$, it is optimal. Moreover, the same recursion implies that the best-case cost is $1$ and the worst-case cost is $3$.
Step 2: optimal policy on noiseless CUBE-NM. If $n=1$, CUBE-NM reduces to the noiseless CUBE dataset (because dummy features are irrelevant with $\sigma=0$), so $\mathbb{E}_{\mathbf{x}}\mathbb{E}_{\pi^{\ast}}[c(\pi^{\ast}[\mathbf{x}])]=2.25$, $\min_{x\in\mathcal{X}}c(\pi^{\ast}[x])=1$, and $\max_{x\in\mathcal{X}}c(\pi^{\ast}[x])=3$.
Assume $n\geq 2$. Querying $f_{1}$ reveals the active block in exactly one query. After that, the remaining task is exactly noiseless CUBE on the active block, which has optimal expected cost $2.25$ by Step 1. Hence the policy “query $f_{1}$ first, then run an optimal noiseless CUBE policy inside the active block” has expected cost $1+2.25$, best case $1+1$, and worst case $1+3$.
This is optimal: any query in an inactive block always returns $0.5$ and provides no information about $y$, so such queries can only increase cost. Since $n\geq 2$, any policy must determine which block is active before it can reliably use informative features, and querying $f_{1}$ does this in one step. Therefore $\mathbb{E}_{\mathbf{x}}\mathbb{E}_{\pi^{\ast}}[c(\pi^{\ast}[\mathbf{x}])]=\mathbf{1}\{n\geq 2\}+2.25$, $\min_{x\in\mathcal{X}}c(\pi^{\ast}[x])=\mathbf{1}\{n\geq 2\}+1$, and $\max_{x\in\mathcal{X}}c(\pi^{\ast}[x])=\mathbf{1}\{n\geq 2\}+3$.
Step 3: myopic CMI policy. Fix $n=n_{c}$. Recall that CUBE-NM has one context feature $f_{1}$ (encoding the active context) and $n$ blocks of 10 features. Denote the block features by $g_{b,i}$ for block $b\in\{1,\dots,n\}$ and within-block index $i\in\{1,\dots,10\}$. In the noiseless setting, if block $b$ is inactive then $\mathbf{x}_{g_{b,i}}=0.5$ deterministically, while in the active block $j$ we observe $\mathbf{x}_{g_{j,i}}=x^{(y)}_{i}$.
Intuition for the behavior of $\pi_{\mathrm{CMI}}$. By construction the context is independent of the label, so $I(\mathbf{y};\mathbf{x}_{f_{1}})=0$ and $\pi_{\mathrm{CMI}}$ does not query $f_{1}$ as long as there exist block features with strictly positive CMI. Initially, the CMI maximizers are the four within-block indices $i\in\{4,5,6,7\}$ in any block. After querying some $g_{b,i}$ and observing $0.5$, block $b$ becomes less likely to be the active block than any untouched block, because the outcome $0.5$ is guaranteed if $b$ is inactive but occurs only for a strict subset of labels if $b$ is active. Consequently, it can be verified (by a direct CMI calculation) that $\pi_{\mathrm{CMI}}$ prefers to query the same within-block index $i$ in an untouched block rather than querying a different feature in the same block. In effect, at a fixed within-block index $i$, $\pi_{\mathrm{CMI}}$ tests blocks one by one until it either observes a non-$0.5$ value (which can only happen in the active block) or it has tested all $n$ blocks and concludes that the active block also has value $0.5$ at index $i$. Once the value at index $i$ in the active block is determined, $\pi_{\mathrm{CMI}}$ moves to the next within-block index that maximizes CMI under the reduced label set, and repeats the same block-testing behavior. The exact comparisons of CMI are straightforward to verify for the noiseless prototypes, so they are omitted for brevity.
Expected cost. The policy $\pi_{\mathrm{CMI}}$ selects, at each step, a feature with maximal CMI $I(\mathbf{y};\mathbf{x}_{a}\mid\mathbf{x}_{S})$, breaking ties uniformly at random. As argued above, its behavior can be viewed as a sequence of stages. In a stage, $\pi_{\mathrm{CMI}}$ fixes a within-block index $i$ (chosen as a CMI maximizer under the current posterior on $\mathbf{y}$) and then queries features of the form $g_{b,i}$ across blocks $b$ until it has determined the value of the active block at index $i$ (either by observing a non-$0.5$ value in some block, or by exhausting all $n$ blocks and concluding that the active block also has value $0.5$). After this, the posterior on $\mathbf{y}$ is reduced, a new maximizer $i$ is selected, and the next stage begins.
Block-testing cost within a stage.
Fix a stage index $i$ and let
$p\;:=\;\Pr\!\big(x^{(\mathbf{y})}_{i}\neq 0.5\,\big|\,\text{current posterior on }\mathbf{y}\big).$
With probability $p$, the active block produces a non-$0.5$ value at index $i$, and $\pi_{\mathrm{CMI}}$ stops the stage as soon as it tests the active block. With uniform random tie-breaking across blocks, the position of the active block among the tested blocks is uniform on $\{1,\dots,n\}$, so the expected number of block tests in this case is $(n+1)/2$. With probability 1p1-p, the active block also has value $0.5$ at index $i$, so every queried block returns $0.5$ and $\pi_{\mathrm{CMI}}$ must test all $n$ blocks to conclude this. Therefore, the expected number of queries spent in this stage is
$q(p)\;:=\;p\cdot\frac{n+1}{2}\;+\;(1-p)\cdot n.$
First-stage tie and two symmetry classes.
Initially, under the uniform prior on $\mathbf{y}\in\{1,\dots,8\}$, the CMI maximizers are exactly the indices $i\in\{4,5,6,7\}$, and the first queried feature is uniformly random among $\{g_{b,i}:b\in\{1,\dots,n\},\,i\in\{4,5,6,7\}\}$. By symmetry, the event $i\in\{4,7\}$ occurs with probability $1/2$ and the event $i\in\{5,6\}$ occurs with probability $1/2$. These two cases lead to different reduced label sets after the first stage, and hence to different expected remaining costs. Let $G(n)$ denote the expected total cost conditional on the first-stage index belonging to $\{4,7\}$, and let $B(n)$ denote the analogous quantity conditional on the first-stage index belonging to $\{5,6\}$.
Case A: first stage uses $i\in\{4,7\}$.
By symmetry it suffices to analyze $i=4$. From the noiseless CUBE prototypes, exactly $3$ of the $8$ labels satisfy $x^{(y)}_{4}\neq 0.5$, so the first stage has $p=3/8$ and expected cost $q(3/8)$. If the active-block value is non-$0.5$, then with probability $2/3$ of that event the observation corresponds to two labels, which requires one additional query to resolve; this contributes $(2/8)\cdot 1$ to the expectation. If instead the active-block value is $0.5$ (probability $5/8$), the posterior reduces to a 5-label set, and the next myopic maximizer is $i=7$. On this 5-label set, exactly $3$ labels satisfy $x^{(y)}_{7}\neq 0.5$, so the second stage has $p=3/5$ and expected block-testing cost $q(3/5)$. If this stage yields a non-$0.5$ value in the active block, then with probability $2/3$ of that event two labels remain and one further query resolves the label; this contributes $(2/5)\cdot 1$ within the 5-label branch. If instead the active-block value is again $0.5$ (probability $2/5$ within the 5-label branch), then exactly two labels remain and the final distinguishing stage has $p=1/2$, contributing $q(1/2)$. Altogether,
$G(n)$ $=q(3/8)+(2/8)\cdot 1+(5/8)\,H(n),$
$H(n)$ $=q(3/5)+(2/5)\cdot 1+(2/5)\,q(1/2),$
which simplifies to $G(n)=(23n+15)/16$.
Case B: first stage uses $i\in\{5,6\}$.
By symmetry it suffices to analyze $i=5$. Again exactly $3$ of the $8$ labels satisfy $x^{(y)}_{5}\neq 0.5$, so the first stage has $p=3/8$ and cost $q(3/8)$, and the non-$0.5$ outcome yields a two-label ambiguity with probability $2/8$, contributing $(2/8)\cdot 1$. If the active-block value is $0.5$ (probability $5/8$), the posterior reduces to a 5-label set and the next myopic maximizer is $i=2$, for which exactly $2$ of the $5$ labels satisfy $x^{(y)}_{2}\neq 0.5$; hence the next stage has $p=2/5$ and cost $q(2/5)$. If the active-block value at $i=2$ is non-$0.5$, then one additional query resolves the remaining two labels; if it is $0.5$ (probability $3/5$ within this branch), then the posterior reduces to a 3-label set. On that 3-label set, the myopic continuation performs a stage with $p=1/3$ (detecting a single label) and, if needed, a final stage with $p=1/2$ (separating the last two labels). Thus
$B(n)$ $=q(3/8)+(2/8)\cdot 1+(5/8)\,M(n),$
$M(n)$ $=q(2/5)+(3/5)\,C(n),$
$C(n)$ $=(1/3)\cdot\frac{n+1}{2}+(2/3)\Big(n+q(1/2)\Big)=\frac{4n+1}{3},$
which simplifies to $B(n)=(29n+11)/16$.
Averaging over the initial tie.
Finally, averaging over the first-stage symmetry classes (each with probability $1/2$) gives
$\mathbb{E}_{\mathbf{x}}\mathbb{E}_{\pi_{\mathrm{CMI}}}\!\big[c(\pi_{\mathrm{CMI}}[\mathbf{x}])\big]=\frac{G(n)+B(n)}{2}=\frac{13(2n+1)}{16}.$
Best-case cost. There exist labels $y$ for which a single informative query in the active block uniquely identifies the label (for example, at index $i=4$ the value 0 occurs for exactly one class). Since $\pi_{\mathrm{CMI}}$ may select such a feature in the active block on the first step, $\min_{x\in\mathcal{X}}c(\pi_{\mathrm{CMI}}[x])=1$.
Worst-case cost. There exist labels $y$ for which three successive myopic stages require exhausting all $n$ blocks before a non-$0.5$ value is observed or the stage is concluded (e.g., the active block yields $0.5$ at each of the three within-block indices selected by the myopic rule). In that event each stage costs $n$ block tests, so $\max_{x\in\mathcal{X}}c(\pi_{\mathrm{CMI}}[x])=3n$. ∎